Grade 9Algebra

(1+2+3+4+5+6+7+8+9+10+11+12+13+14+15+16+17+18+19+20+21+22+23+24+25+26+27+28+29+30+31+32+33+34+35+36+37+38+39+40+41+42+43+44+45+46+47+48+49+50)2. solve

tarang pandey
6 Years agoGrade 9

2 Answers

Arun
6 Years ago
Dear student
 
first calculate sum of 1 to 50 numbers = 50 * 51/2 = 1275
 
Now we square
 
(1275)2 = 1275 * 1275 = 1625625
 
Hope it helps
 
Regards
Arun
Vikas TU
6 Years ago
Dear student 
The sum of n terms of natural number = n(n+1) /2 
So the number is upto n , 
Sum = 50 * 51 /2 
= 1275 
Now you can take square of this 
(1275)^2 will be the ans ,
So, (1+2+3+4+5+6+7+8+9+10+11+12+13+14+15+16+17+18+19+20+21+22+23+24+25+26+27+28+29+30+31+32+33+34+35+36+37+38+39+40+41+42+43+44+45+46+47+48+49+50)2. = 1625625
Hope this helps 
Good Luck